OpenStudy (anonymous):

find limit as x approaches 0 from the positive side. f(x)= x^sinx , answer is one

OpenStudy (anonymous):

f(x) = x^sinx = e^(sinx x lnx) as x approaches 0, sinx x lnx = 0 so, f(x) = e^0 = 1